OpenStudy (vuriffy):

Integrate:

OpenStudy (vuriffy):

That's an a on the limit, by the way. This one scares me, lol.

OpenStudy (karim728):

is dat 1/x times x?

hartnn (hartnn):

what is this term? \(3 e^{(1/x) \times x}\)

OpenStudy (vuriffy):

Would that become 3e^1 ?

hartnn (hartnn):

1/x * x is just 1, yeah

OpenStudy (karim728):

yeap then break it into 2 integrals

hartnn (hartnn):

just 3e+1 then? its a constant, just bring it out!

OpenStudy (vuriffy):

That's it...?

OpenStudy (karim728):

now integrate

OpenStudy (vuriffy):

I saw the worksheet like this I have, but it didn't make much sense, is there like a constant rule?

hartnn (hartnn):

(3e+1) \(\int_0^a dx = (3e+1)a\)

hartnn (hartnn):

constant can be brought out of the integral \(\int af(x)dx = a\int f(x) dx\)

hartnn (hartnn):

karim, the x in the exponent cancelled, right?

hartnn (hartnn):

vuriffy, can u post the screenshot of the question? just to be sure

OpenStudy (vuriffy):

It's on my phone, I can write it exactly how it is in a drawing.

OpenStudy (karim728):

ooh yeh they canceled

OpenStudy (vuriffy):

Oh crap, I wrote it wrong.

OpenStudy (vuriffy):

 Some reason it was blurred, but now the picture like became fully clear.